OpenStudy (yumeybitaa):

A square has a side of 6.25 feet. What is the area in square feet? Round to the correct significant digits

OpenStudy (crashonce):

whats the area of a square formula?

OpenStudy (crashonce):

it is area=one side squared so the area will be 6.25^2 = 39.0625 square feet

OpenStudy (anonymous):

see the formula for the area of square=lenght into width so for sqr. length=width and here l=w=6.25 so area=6.25*6.25= ??? answer.